1. Solve the given differential equation using an appropriate substitution. The DE is a Bernoulli equation,

dy/dx = y(xy^6 - 1)

2. Solve the given differential equation by using an appropriate substitution. The DE is a Bernoulli equation.

x dy/dx + y = 1/y^2
1
Expert's answer
2020-07-20T18:00:59-0400
"Solution"

1. To solve the DE Bernoulli equation,

"dy\/dx = y(xy^6 - 1)"


The standard form is

"\frac{dy}{dx}+P(x)y=f(x)y^n"


"Where" "n" "is" "not" "equal" "to" "zero" "or" "one"


"dy\/dx = xy^7 - y"

"dy\/dx+y = xy^7" "-------------->(1)"


"n=7"

"U=y^{1-n}=y^{1-7}=y^{-6}"

"U=y^{-6}"

"U^{-\frac{1}{6}}=(y^{-6})^{-\frac{1}{6}}" "\implies" "y=U^{-\frac{1}{6}}"

"dy/dx =- \frac{1}{6}U^{- \frac{7}{6}}*\frac{du}{dx}"


"\frac{1}{6}U^{- \frac{7}{6}}\frac{du}{dx}+U^{-\frac{1}{6}}=x U^{- \frac{7}{6}}----------->(2)"


Multiply by "( \frac{1}{6}U^{- \frac{7}{6}})"


"\frac{du}{dx}-6U^1=-6x"


"y(x)=\epsilon^{\int \rho (x)dx}=\epsilon^{\int -6dx}= \epsilon^{-6x}"

"\implies y(x)=\epsilon^{-6x}"


"\epsilon^{-6x} \frac{du}{dx}-6 \epsilon^{-6x} U=-6x\epsilon^{-6x}"


"\frac{d}{dx}[\epsilon^{-6x} U]=-6x\epsilon^{-6x}"


Integration:

"\epsilon^{-6x} U=\int -6x\epsilon^{-6x} dx"


"\epsilon^{-6x} U=x\epsilon^{-6x} +\frac{1}{6}\epsilon^{-6x} +C"


"\implies U=x+\frac{1}{6}+\frac{C}{\epsilon^{-6x} }"

"y^{-6}=x+\frac{1}{6}+C\epsilon^{6x} ------------->Answer"




2. To solve the DE Bernoulli equation.


"x \frac{dy}{dx} + y = \frac{1}{y^2}"


The standard form is

"\frac{dy}{dx}+P(x)y=f(x)y^n"


Where n is not equal to zero or one


"\frac{dy}{dx}+\frac{1}{x}y=\frac{1}{x}y^{-2}------------>(1)"


"n=-2"

"U=y^{1-n}=y^{1-(-2)}=y^3"

"U=y^{3}"

"\implies y=U^{\frac{1}{3}}"


"\frac{dy}{dx}=\frac{1}{3}U^{-\frac{2}{3}}\frac{du}{dx}"


"\frac{1}{3}U^{-\frac{2}{3}}\frac{du}{dx}+\frac{1}{x}U^{\frac{1}{3}}=\frac{1}{x}U^{-\frac{2}{3}}------->(2)"


Multiply by "(3U^{-\frac{2}{3}})"


"\frac{du}{dx}+\frac{3}{x}U^1=\frac{3}{x}"


"y(x)=\epsilon^{\int \rho (x)dx}=\epsilon^{\int \frac{3}{x}dx}= \epsilon^{3ln|x|}=|x^3|=x^3"

"\implies y(x)=x^3"


"x^3\frac{du}{dx}+3x^2U=3x^2"


"\frac{d}{dx}[x^3U]=3x^2"


"x^3U=\frac{3x^2}{3}+C"


"U=1+\frac{C}{x^3}"

"y^3=1+\frac{C}{x^x} \implies y=(1+\frac{C}{x^x})^{\frac{1}{3}}------->Answer"